Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Entry  Fri Jun 28 22:49:39 2019, Dirk Hoffmann, Telescope Operations, General, Telescope stuck,  
    Reply  Fri Jun 28 23:06:36 2019, Dirk Hoffmann, Telescope Operations, General, Telescope stuck,  
Message ID: 258     Entry time: Fri Jun 28 23:06:36 2019     In reply to: 257
Author: Dirk Hoffmann 
Type: Telescope Operations 
Category: General 
Subject: Telescope stuck 
Status:  

Ino explained: MOS was totally desynchronised with PLC. (It was assuming that the telescope was parked.) He parked the telescope manually. Thanks!
Experts will investigate on Monday. But the shift crews can continue to use the  Drive Control and tracking.

Quote:

After tracking Mk421, we abort current action (acknowledged) and try to go to a new source: RA=308.1, Dec=40.96. Drive Control refuses these parameters ("check your parameters").

General Abort without effect for 2 minutes. Called expert (Ino). He will look into it.

 

ELOG V3.1.5-fc6679b